# Relationship between the ray variability and the pc-scale jet in the blazar 3C 454.3

### v1

@@ Abstract
Context. 3C 454.3 is a flat spectrum radio quasar (FSRQ) known for its high variability across the electromagnetic spectrum, showing structural and flux variability in its pc-scale jet, and correlated variability among frequency bands.
Aims. This study aims to identify the structure, dynamics, and radiative processes common to the innermost regions of the blazar 3C 454.3. We investigate whether any jet component can be associated with ray emission and variability. Additionally, we compare the flux variability of the pc-scale jet components found in the VLBA observations to the variability in the ray emission.
Methods. We analyze the relationship between the variable ray emission and pc-scale jet properties in 3C 454.3 by combining ray data spanning twelve years with contemporaneous VLBA multi-epoch images at 15 and 43 GHz. Spearmanās rank correlation tests are conducted to determine if the flux variability of any jet component is associated with ray variability.
Results. Core emission at 43 and 15 GHz strongly correlates with ray emission. The 43 GHz core (Q0) contributes around 37 of the observed ray variability, while the 15 GHz core (K0) accounts for 30. A quasi-stationary component at 43 GHz, at a projected distance of 4.6 pc, correlates with the ray flux, accounting for 20 of its emission between 2016 and 2021. We found a mobile component (Q3 between 2010.18 and 2011.16) at 43 GHz with a projected distance between 0.8 and 2.3 pc and apparent velocity of c, accounting for approximately 28 of the ray emission. The observed simultaneous variability in emission regions beyond the central parsec strongly suggests synchrotron self-Compton (SSC) as the primary mechanism for ray production in these regions.
Conclusions. Our findings demonstrate the existence of multiple ray emission regions within the blazar jet but also suggest that some of these regions are non-stationary over time. Additionally, our study pinpoints the exact locations of these emission regions within the blazar itself. These results are valuable for both theoretical models and for gaining a deeper understanding of the complex nature of blazars.
@@
Blazars (Angel & Stockman, 1980) are among the least common class of active galactic nuclei (AGN) whose emission is dominated by radiation from nonāthermal plasma moving at relativistic speed in a direction close to the line of sight. Given the jet small viewing angle (e.g., Hovatta & Lindfors, 2019) blazar emission shows relativistic effects such as beaming and Doppler boosting, which causes the jet components to move with apparent superluminal speeds up to tens of the speed of light (e.g., Jorstad etĀ al., 2017). Although, blazars make up the majority of discrete ray sources seen by the Fermi Large Area Telescope (LAT, Abdollahi etĀ al., 2022), what are the mechanisms of ray production, where they occur or if there can be different mechanisms operating even in the same object, remain open questions. The spectral energy distribution (SED) of blazars shows two peaks (Ghisellini, 1989; Dermer & Schlickeiser, 1993; Sikora etĀ al., 1994): the first one due to synchrotron emission, and the second one, at higher energies, generally attributed to inverse Compton (IC) scattering. The origin of the IC seed photons has been a topic of years-long debate: these could come from the accretion disk emission, or from the broad line region (BLR, Sikora etĀ al., 1994) or from the dusty torus (Sikora etĀ al., 2008), a process known as external inverse Compton (EC) scattering. Alternatively, the seed photons can originate from synchrotron emission produced in the jet itself, a mechanism called synchrotron self-Compton (SSC, Maraschi etĀ al., 1992; Bloom & Marscher, 1996). Also, hadronic mechanisms for producing the second peak have been proposed (e.g., Mannheim, 1993).
@@
Correlated variability between different wavelengths can help us locate the ray emission region within the jet. However, Very Long Baseline Interferometry (VLBI) offers an even more powerful tool. VLBI observations at millimeter and radio wavelengths achieve the highest spatial resolution, allowing us to pinpoint the exact location, or āblazar region,ā where rays are produced (see review by Boccardi etĀ al., 2017). Several studies have provided valuable insights into the location and physical processes responsible for ray emission in AGN. For instance, (Jorstad etĀ al., 2001; LƤhteenmƤki & Valtaoja, 2003; León-Tavares etĀ al., 2011; Fuhrmann etĀ al., 2014) suggest that radio emission often occurs at the same time or after rays. On the other hand, ray emission has been detected from the blazar 3C 84 in multiple locations (Hodgson etĀ al., 2018; Paraschos etĀ al., 2023), as well as from the blazars 1633+382 and 3C 345 (Ramakrishnan etĀ al., 2015). León-Tavares etĀ al. (2013) linked variations in the Mg II 2800 emission line with radio emission from superluminal jet material in 3C 454.3. This finding suggests the presence of broad emission line material well beyond the inner parsec, close to the radio core. Similarly, Isler etĀ al. (2013) provided evidence for a emitting region interior to the Broad Line Region (BLR) during flares in 2009 and 2010 for the same object. Recent VLBI observations have revealed even more complexity. PatiƱo-Ćlvarez etĀ al. (2019) studied 3C 279 and found, for the first time, evidence for a non-stationary ray emission region located about 42 pc from the core. Their results suggest that the ray emission may originate from different places within the jet. Likewise, Chavushyan etĀ al. (2020) found evidence of BLR clouds at about 25 pc from the supermassive black hole in CTA 102. Further supporting the dynamic nature of blazar jets, Traianou etĀ al. (2024) analyzed VLBI data of 3C 454.3 at high frequencies (43 GHz and 86 GHz) between 2013 and 2017. They observed four jet knots that seemingly disappeared upon reaching a quasi-stationary component at about 0.6 mas. The authors propose that this could be due to the jet folding back onto itself, redirecting the plasma toward our line of sight.
@@
3C 454.3 is known as one of the brightest sources in the ray sky. Between 2009 and 2010, it presented a series of important ray flares (Ackermann et al., 2010; Vercellone et al., 2011). The late-2010 flare stands out as the brightest ever recorded at the time, reaching a peak flux of photons cm-2 s-1 (Abdo et al., 2011). This translates to an isotropic luminosity of approximately erg (Sahakyan, 2021), solidifying its place among the brightest ray sources known. Extensive research has confirmed a strong correlation between the ray variability and radio frequencies, as well as other wavelengths (e.g., Jorstad et al., 2010; León-Tavares et al., 2013; Liodakis et al., 2020; Amaya-AlmazÔn et al., 2021). The passage of a plasma component down the jet is believed to be a primary driver for most ray flares (e.g., Marscher et al., 2010; Jorstad et al., 2013, 2017; Weaver et al., 2019). In some cases, the flares coincide with the interaction of superluminal components within the jet, such as their passage through a quasi-stationary feature or the core (e.g., Agudo et al., 2011). For instance, Jorstad et al. (2013) proposed that the 2009-2010 flares in 3C 454.3 might be linked to the emergence and propagation of two knots within the jet, designated K09 and K10, as identified through analysis of monthly Very Long Baseline Array (VLBA) images. Similarly, Weaver et al. (2019) identified K16, a component ejected around 4 months before the ray flare of 2016, which was also observed in optical. They suggest that K16 is likely responsible for the 2016 outburst when its lagging end moved past a standing shock in the core. More recently Liodakis et al. (2020) reported that the ray flare of 2014 might be the result of a traveling shock, the component K14, when it crossed a standing re-collimation shock observed as the radio core. These events are illustrated in Fig. 1 with vertical lines.
@@
This study investigates the blazar 3C 454.3 using data from the Very Long Baseline Array (VLBA) at radio frequencies of 43 GHz and 15 GHz. We analyze these observations alongside ray flux variations to explore the jet structure, dynamics, emission mechanisms, and the location of ray production regions. Section 2 details the observational data and the analysis methods employed. Section 3 presents our results and discusses the jet physical properties. Finally, Section 4 summarizes our key findings. The cosmological parameters adopted throughout this paper are km , , (Planck Collaboration etĀ al., 2014). At 3C 454.3 redshift (z = 0.859), the distance to the object is estimated to be 5.489 Gigaparsecs (Gpc). Given this distance, one milliarcsecond (mas) on the sky corresponds to a physical size of 7.7 parsecs (pc).
@@ 2 Observational data and analysis
This section describes the long-term radio and ray datasets used in our analysis of 3C 454.3, along with the methods we employed to analyze them. All the corresponding light curves are presented in Fig. 1.
2.1 Gamma rays.
We obtained ray data for 3C 454.3 from the Fermi Large Area Telescope (Fermi-LAT) public archive, covering energies between 0.1 and 300 GeV (Abdo etĀ al., 2010). To process and analyze these data, we used the Fermitools software (version 2.0.8). During the analysis, we considered all ray sources within a 15-degree radius of 3C 454.3, as listed in the 4FGL-3DR catalog (Abdollahi etĀ al., 2022). We allowed the model to freely adjust the intensity of these nearby sources. For sources even closer (within a 5-degree radius), we allowed the model to adjust all their spectral parameters in addition to intensity. This helps to account for potential contamination from other ray emitters in the region closer to our source of interest. For our main analysis, we focused on ray flux measurements centered on the specific dates when VLBA observations were made. Additionally, to illustrate the variability in the ray emission, we created a weekly light curve spanning from August 2008 to late 2020. This light curve is shown in Fig. 1a, but it is important to note that this is for demonstration purposes only and not used in the analysis.
@@ 2.2 VLBA observations at 43 and 15 GHz
We used 137 VLBA maps at 43 GHz (7 mm wavelength) from the Boston University VLBA Blazar monitoring program (VLBA-BU-BLAZAR) and its successor Blazars Entering the Astrophysical Multi-Messenger Era (BEAM-ME, Jorstad etĀ al., 2017), to study the pc-scale jet kinematics and structural evolution during the same period as the ray light curve (2008 - 2020)111http://www.bu.edu/blazars/BEAM-ME.html. The program details, observations, and data reduction are described in Jorstad etĀ al. (2017). VLBA observations at 43 GHz typically achieve angular resolutions on the order of microarcseconds (as) and sensitivities around a few tens of Jy per beam. In the case of 3C 454.3 the angular resolution is on the order of 0.1 mas.
Additionally, we included 54 images at 15 GHz from Monitoring of Jets in Active Galactic nuclei with VLBA Experiments (MOJAVE, Lister etĀ al. 2009)222https://www.cv.nrao.edu/MOJAVE/project.html. The images were taken from mid-2008 to late 2021. However, data from 2019 was excluded due to a known systematic effect causing a 10% decrease in the flux density starting from the first half of 2019333See also: https://science.nrao.edu/enews/14.4/index.shtml#vlba_flux.. VLBA observations at 15 GHz can achieve sensitivities in the range of tens to hundreds of Jy/beam and angular resolution on the order of tens to hundreds of as. For 3C 454.3 the angular resolution varies depending on the number of antennas available and configuration but has a typical value of 0.3 mas.
@@ 2.3 OVRO 15 GHz observations
To complement the high-resolution VLBA data, we also examined the total flux density of 3C 454.3 at 15 GHz using data from the Owens Valley Radio Observatory (OVRO) blazar monitoring program (Richards etĀ al., 2011)444https://sites.astro.caltech.edu/ovroblazars/. This data provides a broader view of the blazar emission compared to the VLBA observations, which focus on jet features. The light curve showing the variations in the total flux density at 15 GHz is included in Fig. 1d.
2.4 Sub-Millimeter Array 230 and 350 GHz.
To complement the radio data, we analyzed the blazar emission at higher frequencies of 230 and 350 GHz (wavelengths of 1.3 and 0.8 millimeters). The data was obtained from the publicly available Sub-Millimeter Array (SMA) archive555http://sma1.sma.hawaii.edu/smaoc.html. The light curve showing the variations in flux density at this frequency is included in Fig. 1b. For details on the observations and data processing, please refer to Gurwell etĀ al. (2007).
2.5 Jet structure model-fitting
To characterize the pc-scale jet structure from the VLBA images we followed the methodology described in PatiƱo-Ćlvarez etĀ al. (2019). We used the Common Astronomy Software Applications (CASA) task IMFIT to model the emission of each component, including the core, using 2D elliptical Gaussian components. A model feature was considered a detection if it equaled or exceeded 6 from the background emission. For both frequencies, we identified the core as the easternmost component, along with other relevant knots in each map. Each component was cross-identified between different epochs by comparing the integrated flux and position relative to the core.
The jet components, or knots, were named according to the frequency at which they were observed. Specifically, Q designates those observed at 43 GHz and K those identified at 15 GHz. Knots are assigned sequential numbers according to the order in which they were observed within the jet; these designations also reflect the distance of the knot from the core, although this is not always the case. Components identified in fewer than 5 epochs were omitted (such as the components Q9 and Q16).
@@ 2.5 Jet structure model-fitting
The observations made using high angular resolution VLBA have revealed that 3C 454.3 can be categorized as core-dominated between 2008 and 2020, with a few exceptional time intervals where the flux is dominated by some other component, displaying an intricate multi-component jet structure. Fig. 2 illustrates the results of modeling using the task IMFIT near the core of 3C 454.3 within a margin of 2 and 10 mas at 43 and 15 GHz, respectively. We also tracked the temporal evolution of the radial separation between the different jet components and the radio core between 2008 and 2020 (see Fig. 3), enabling the calculation of their proper motions and apparent velocities (see Fig. 4). The light curves of all identified jet components in 3C 454.3 at 43 and 15 GHz are shown in Fig. 5.

### v2

# Relationship between the -ray variability and the parsec-scale jet in the blazar 3C 454.3

@@ Abstract
Context. The flat spectrum radio quasar 3C 454.3 is known for its high variability across the electromagnetic spectrum, showing structural and flux variability in its parsec-scale jet and correlated variability among frequency bands.
Aims. This study aims to identify the structure, dynamics, and radiative processes common to the innermost regions of the blazar 3C 454.3. We investigate whether any jet component can be associated with -ray emission and variability. Additionally, we compare the flux variability of the parsec-scale jet components found in the VLBA observations to the variability in the -ray emission.
Methods. We analyzed the relationship between the variable -ray emission and parsec-scale jet properties in 3C 454.3 by combining -ray data spanning twelve years with contemporaneous VLBA multi-epoch images at 15 and 43 GHz. We conducted Spearmanās rank correlation tests to determine if the flux variability of any jet component is associated with -ray variability.
Results. The core emission at 43 and 15 GHz strongly correlates with -ray emission. The 43 GHz core (Q0) contributes around 37 of the observed -ray variability, while the 15 GHz core (K0) accounts for 30. A quasi-stationary component at 43 GHz, at a projected distance of 4.6 pc, correlates with the -ray flux, accounting for 20 of its emission between 2016 and 2021. We found a mobile component (Q3 between 2010.18 and 2011.16) at 43 GHz with a projected distance between 0.8 and 2.3 pc and an apparent velocity of c that accounts for approximately 28 of the -ray emission. The observed simultaneous variability in emission regions beyond the central parsec strongly suggests synchrotron self-Compton as the primary mechanism for -ray production in these regions.
Conclusions. Our findings demonstrate the existence of multiple -ray emission regions within the blazar jet but also suggest that some of these regions are non-stationary over time. Additionally, our study pinpoints the exact locations of these emission regions within the blazar itself. These results are valuable for theoretical models and for gaining a deeper understanding of the complex nature of blazars.
@@
Blazars (Angel & Stockman, 1980) are among the least common class of active galactic nuclei (AGN), and their emission is dominated by radiation from nonāthermal plasma moving at a relativistic speed in a direction close to the line of sight. Given the small viewing angle of the jet (e.g., Hovatta & Lindfors, 2019), blazar emission shows relativistic effects such as beaming and Doppler boosting that cause the jet components to move with apparent superluminal speeds of up to ten times the speed of light (e.g., Jorstad etĀ al., 2017). Although blazars make up the majority of discrete -ray sources seen by the Fermi Large Area Telescope (Abdollahi etĀ al., 2022), such questions as what the mechanisms of -ray production are, where they occur, and if there can be different mechanisms operating even in the same object remain open. The spectral energy distribution of blazars shows two peaks (Ghisellini, 1989; Dermer & Schlickeiser, 1993; Sikora etĀ al., 1994). The first one is due to synchrotron emission, and the second one, at higher energies, is generally attributed to inverse Compton (IC) scattering. The origin of the IC seed photons has been a topic of years-long debate. They could come from the accretion disk emission or from the broad-line region (BLR; Sikora etĀ al., 1994) or from the dusty torus (Sikora etĀ al., 2008), a process known as external IC (EIC) scattering. Alternatively, the seed photons can originate from synchrotron emission produced in the jet itself, a mechanism called synchrotron self-Compton (SSC, Maraschi etĀ al., 1992; Bloom & Marscher, 1996). Hadronic mechanisms for producing the second peak have also been proposed (e.g., Mannheim, 1993).
@@
The correlated variability between different wavelengths can help locate the -ray emission region within a jet. However, Very Long Baseline Interferometry (VLBI) offers an even more powerful tool. VLBI observations at millimeter and radio wavelengths achieve the highest spatial resolution, thus allowing the exact location, or āblazar region,ā where -rays are produced to be identified (see review by Boccardi etĀ al., 2017). Several studies have provided valuable insights into the location and physical processes responsible for -ray emission in AGN. For instance, (Jorstad etĀ al., 2001; LƤhteenmƤki & Valtaoja, 2003; León-Tavares etĀ al., 2011; Fuhrmann etĀ al., 2014) suggest that radio emission often occurs at the same time or after -rays. On the other hand, -ray emission has been detected from the blazar 3C 84 in multiple locations (Hodgson etĀ al., 2018; Paraschos etĀ al., 2023) as well as from the blazars 1633+382 and 3C 345 (Ramakrishnan etĀ al., 2015). León-Tavares etĀ al. (2013) linked variations in the Mg II 2800 emission line with radio emission from superluminal jet material in 3C 454.3. This finding suggests the presence of broad emission line material well beyond the inner parsec close to the radio core. Similarly, Isler etĀ al. (2013) provided evidence for a emitting region interior to the broad-line region (BLR) during flares in 2009 and 2010 for the same object. Recent VLBI observations have revealed even more complexity. PatiƱo-Ćlvarez etĀ al. (2019) studied 3C 279 and found, for the first time, evidence for a non-stationary -ray emission region located about 42 pc from the core. Their results suggest that the -ray emission may originate from different places within the jet. Likewise, Chavushyan etĀ al. (2020) found evidence of BLR clouds at about 25 pc from the supermassive black hole in CTA 102. Further supporting the dynamic nature of blazar jets, Traianou etĀ al. (2024) analyzed VLBI data of 3C 454.3 at high frequencies (43 GHz and 86 GHz) between 2013 and 2017. They observed four jet knots that seemingly disappeared upon reaching a quasi-stationary component at about 0.6 mas. The authors propose that this could be due to the jet folding back onto itself and thus redirecting the plasma toward our line of sight.
@@
Notably, 3C 454.3 is known as one of the brightest sources in the -ray sky. Between 2009 and 2010, it presented a series of important -ray flares (Ackermann et al., 2010; Vercellone et al., 2011). The late-2010 flare stands out as the brightest ever recorded at the time, reaching a peak flux of photons cm-2 s-1 (Abdo et al., 2011). This translates to an isotropic luminosity of approximately erg (Sahakyan, 2021), solidifying its place among the brightest -ray sources known. Extensive research has confirmed a strong correlation between the -ray variability and radio frequencies, as well as other wavelengths (e.g., Jorstad et al., 2010; León-Tavares et al., 2013; Liodakis et al., 2020; Amaya-AlmazÔn et al., 2021). The passage of a plasma component down the jet is believed to be a primary driver for most -ray flares (e.g., Marscher et al., 2010; Jorstad et al., 2013, 2017; Weaver et al., 2019). In some cases, the flares coincide with the interaction of superluminal components within the jet, such as their passage through a quasi-stationary feature or the core (e.g., Agudo et al., 2011). For instance, Jorstad et al. (2013) proposed that the 2009-2010 flares in 3C 454.3 might be linked to the emergence and propagation of two knots within the jet, designated as K09 and K10 and identified through analysis of monthly Very Long Baseline Array (VLBA) images. Similarly, Weaver et al. (2019) identified K16, a component ejected around four months before the -ray flare of 2016, which was also observed in optical. They suggest that K16 is likely responsible for the 2016 outburst when its lagging end moved past a standing shock in the core. More recently Liodakis et al. (2020) reported that the -ray flare of 2014 might be the result of a traveling shock, the component K14, when it crossed a standing re-collimation shock observed as the radio core. These events are illustrated in Fig. 1 with vertical lines.
@@
This study investigates the blazar 3C 454.3 using data from VLBA at radio frequencies of 43 GHz and 15 GHz. We analyze these observations alongside -ray flux variations in order to explore the jet structure, dynamics, emission mechanisms, and the location of -ray production regions. Section 2 details the observational data and the analysis methods employed. Section 3 presents our results and discusses the jetās physical properties. Finally, Sect. 4 summarizes our key findings. The cosmological parameters adopted throughout this paper are km , , and (Planck Collaboration etĀ al., 2014). At a redshift of z = 0.859, the distance to 3C 454.3 is estimated to be 5.489 Gpc. Given this distance, 1 mas on the sky corresponds to a physical size of 7.7 pc.
@@ 2 Observational data and analysis
This section describes the long-term radio and -ray datasets used in our analysis of 3C 454.3 as well as the methods we employed to analyze them. All the corresponding light curves are presented in Fig. 1.
2.1 Gamma rays
We obtained -ray data for 3C 454.3 from the Fermi Large Area Telescope public archive, covering energies between 0.1 and 300 GeV (Abdo etĀ al., 2010). To process and analyze these data, we used the Fermitools software (version 2.0.8). During the analysis, we considered all -ray sources within a 15-degree radius of 3C 454.3, as listed in the 4FGL-3DR catalog (Abdollahi etĀ al., 2022). We allowed the model to freely adjust the intensity of these nearby sources. For sources even closer (within a 5-degree radius), we allowed the model to adjust all their spectral parameters in addition to the intensity. This helped account for potential contamination from other -ray emitters in the region closer to our source of interest. For our main analysis, we focused on -ray flux measurements centered on the specific dates when VLBA observations were made. Additionally, to illustrate the variability in the -ray emission, we created a weekly light curve spanning from August 2008 to late 2020. This light curve is shown in Fig. 1a, but it is important to note that this light curve is for demonstration purposes only and that it was not used in the analysis.
@@ 2.2 VLBA observations at 43 and 15 GHz
We used 137 VLBA maps at 43 GHz (7 mm wavelength) from the Boston University VLBA Blazar monitoring program (VLBA-BU-BLAZAR) and its successor, Blazars Entering the Astrophysical Multi-Messenger Era (BEAM-ME; Jorstad etĀ al., 2017), to study the parsec-scale jet kinematics and structural evolution during the same period as the -ray light curve (2008 - 2020).111http://www.bu.edu/blazars/BEAM-ME.html The program details, observations, and data reduction are described in Jorstad etĀ al. (2017). VLBA observations at 43 GHz typically achieve angular resolutions on the order of microarcseconds (as) and sensitivities around a few tens of Jy per beam. In the case of 3C 454.3, the angular resolution is on the order of 0.1 mas.
Additionally, we included 54 images at 15 GHz from the program Monitoring of Jets in Active Galactic nuclei with VLBA Experiments (MOJAVE; Lister etĀ al. 2009).222https://www.cv.nrao.edu/MOJAVE/project.html The images were taken from mid-2008 to late 2021. However, data from 2019 was excluded due to a known systematic effect causing a 10% decrease in the flux density starting from the first half of 2019.333See also: https://science.nrao.edu/enews/14.4/index.shtml#vlba_flux. VLBA observations at 15 GHz can achieve sensitivities in the range of tens to hundreds of Jy/beam and angular resolution on the order of tens to hundreds of as. For 3C 454.3 the angular resolution varies depending on the number of available antennas and their configuration, but it has a typical value of 0.3 mas.
@@ 2.3 OVRO 15 GHz observations
To complement the high-resolution VLBA data, we also examined the total flux density of 3C 454.3 at 15 GHz using data from the Owens Valley Radio Observatory (OVRO) blazar monitoring program (Richards etĀ al., 2011).444https://sites.astro.caltech.edu/ovroblazars/ This data provides a broader view of the blazar emission compared to the VLBA observations, which focus on jet features. The light curve showing the variations in the total flux density at 15 GHz is included in Fig. 1d.
2.4 Sub-Millimeter Array 230 and 350 GHz
To complement the radio data, we analyzed the blazar emission at higher frequencies of 230 and 350 GHz (wavelengths of 1.3 and 0.8 millimeters). The data was obtained from the publicly available Sub-Millimeter Array archive.555http://sma1.sma.hawaii.edu/smaoc.html The light curve showing the variations in flux density at this frequency is included in Fig. 1b. For details on the observations and data processing, we refer the reader to Gurwell etĀ al. (2007).
2.5 Jet structure model fitting
To characterize the parsec-scale jet structure from the VLBA images, we followed the methodology described in PatiƱo-Ćlvarez etĀ al. (2019). We used the Common Astronomy Software Applications (CASA) task IMFIT to model the emission of each component, including the core, using 2D elliptical Gaussian components. A model feature was considered a detection if it equaled or exceeded 6 from the background emission. For both frequencies, we identified the core as the easternmost component along with other relevant knots in each map. Each component was cross-identified between different epochs by comparing the integrated flux and position relative to the core.
The jet components, or knots, were named according to the frequency at which they were observed. Specifically, Q designates those observed at 43 GHz and K indicates those identified at 15 GHz. Knots were assigned sequential numbers according to the order in which they were observed within the jet. These designations also reflect the distance of the knot from the core, although this is not always the case. Components identified in fewer than five epochs were omitted (such as the components Q9 and Q16).
@@ 2.5 Jet structure model fitting
The observations made using high angular resolution VLBA revealed that 3C 454.3 can be categorized as core-dominated between 2008 and 2020, though with a few exceptional time intervals where the flux is dominated by some other component and displays an intricate multi-component jet structure. Figure 2 illustrates the results of modeling using the task IMFIT near the core of 3C 454.3 within a margin of 2 and 10 mas at 43 and 15 GHz, respectively. We also tracked the temporal evolution of the radial separation between the different jet components and the radio core between 2008 and 2020 (see Fig. 3), enabling the calculation of their proper motions and apparent velocities (see Fig. 4). The light curves of all the identified jet components in 3C 454.3 at 43 and 15 GHz are shown in Fig. 5.